OpenStudy (anonymous):

Which of the following fractions can be used in the conversion of 25 m3 to the unit cm3?

OpenStudy (debbieg):

All you do to convert units is to muliply by the correct CONVERSION FACTOR. The conversion factor must have something in the num'r that is = to what's in the den'r, but just different units. What goes in the num'r and what in the den'r, depends on what units you are STARTING WITH and what units you want to END WITH. So if I have, for example, hours, and I want minutes, I need to use the fact that 1 hour = 60 minutes (see, they are equal?). So my conversion factor is EITHER \(\dfrac{ 60~\min }{ 1~hour }\) OR \(\dfrac{ 1~hour }{ 60~min }\) Which I use depend on whether I'm converting from hours to minutes (so I want HOURS to cancel, thus I need HOURS in the den'r (bottom) and minutes in the num'r (top)). Whatever goes in the TOP is what you'll be LEFT WITH. Whatever goes in the BOTTOM is what will CANCEL. So if I'm going FROM hours, TO minutes: How many minutes are in 5.2 hours? \(5.2~hour\cdot\dfrac{ 60~\min }{ 1~hour }=5.2~\cancel{hour}\cdot\dfrac{ 60~\min }{ 1~\cancel{hour} }=(5.2)\cdot 60~\min=312 \) minutes Now just apply that same method to your conversion problem.